Solve the system. Enter answer as an ordered pair (t,v).
1/2t - 1/5v = 5/2
2/3t + 1/4v = 7/12

Answer by stanbon(75887)

1/2t - 1/5v = 5/2
2/3t + 1/4v = 7/12
---
Get rid of those fractions:
5t - 2v = 25
8t + 3v = 7
----
Multiply 1st by 3
Multiply 2nd by 2
----
15t - 6v = 75
16t + 6v = 14
-----
Add and solve for "t":
31t = 89
t = 89/31
----
solve for "v":
5t - 2v = 25
5*(89/31) - 2v = 25
2v = 445/31 - 25
---
2v = -330/31
v = -165/31
==================
Cheers,
Stan H.
========================